OpenStudy (anonymous):

lim x=0 ((1/2+x) - (1/2))/x

OpenStudy (anonymous):

It grows with out bound as x tends to zero

OpenStudy (anonymous):

@Jack111 o?$$\lim_{x\to0}\frac{1/2+x-1/2}x=\lim_{x\to0}\frac{x}x=\lim_{x\to0}1=1$$

OpenStudy (anonymous):

@oldrin.bataku I am not sure how you are interpreting the tex, it isn't formatted properly

OpenStudy (anonymous):

oh now I see, not watching my parenthesis lol

OpenStudy (cggurumanjunath):

1